# Assessing treatment efficacy for interval-censored endpoints using multistate semi-Markov models fit to multiple data streams

### v2

@@ 4.2 Model for clinical and immunological progression
We fit $6$ models, summarized in Table 4, that were distinguished by the parameterization of their baseline transition intensities, i.e., $\eta_{k\ell}(\tau_{n},\boldsymbol{\gamma}_{k\ell})$ in equation (2). The most flexible model used piecewise linear baseline intensities with 3 interior knots at 3.5, 10.5, and 17.5 days for the transitions from naïve to PCR+ and from PCR+ to symptomatic, and 1 interior knot at 1 week for each of the PCR+ to PCR- transitions. A Markov model with time-homogeneous exponential intensities was the most rigid fully-parametric model. All intensities were adjusted for mAb assignment via a proportional intensity parameterization.
@@ 4.2.1 Seroconversion and asymptomatic infection
Let $C_{A}$, $C_{S}$ and $C=C_{A}+C_{S}$ denote the number of asymptomatic, symptomatic, and total seroconversions observed in an arm. Also let $\widehat{I}_{A}$, $\widehat{I}_{S}$, and $\widehat{I}$ be the analogous number of infections estimated by the model. The estimated probabilities of seroconversion are $\widehat{V}_{A}=C_{A}/\widehat{I}_{A}$, $\widehat{V}_{S}=C_{S}/\widehat{I}_{S}$, and $\widehat{V}=C/\widehat{I}$. Note that we use the estimated probability of asyptomatic infection since otherwise $\widehat{I}_{S}+\widehat{I}_{A}$ may not equal $\widehat{I}$, which would not be coherent under the model. Our use of the bootstrap is motivated by the observation that model-based asymptotic confidence intervals will understate the uncertainty in seroconversion following asymptomatic infection by virtue of conditioning on the number of asymptomatic seroconversion events.
@@ 4.3 Results
Table 4 compares the fitted models in terms of AIC. To showcase the improvement of the semi-Markov models over the Markov model, the difference in log-likelihood with the Markov model is displayed. The semi-Markov models offer a gain in log-likelihood between 78.7 and 101 units and the trade-off at the cost of additional model parameters is worthwhile, as reflected in the smaller AIC for more flexible models. The preferred model by AIC was parameterized with degree 1 B-splines with 2 interior knots for the $1\rightarrow 2$ and $2\rightarrow 4$ transitions, and 1 interior knot for the $2\rightarrow 3$ and $4\rightarrow 5$ transitions. The Markov model exponential baseline intensities had, by far, the worst AIC, while the model with Weibull intensities was worst among the semi-Markov models.
Parameterization of baseline intensities Results
Weibull Weibull 78.7 -133.3 (0.95)
Exponential Exponential 0.0 16.0 (0.0)
Table 4: Model comparison results. The degree and interior knots of the B-Spline intensities are $\kappa$ and $\xi$, respectively. The quantity $\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y})$ indicates the absolute gain in log-likelihood over the Markov model with exponential transition intensities (larger is better). $\Delta$AIC is computed as $-2(\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y}))+2p$ (smaller is better), where $p$ is the number of model parameters. The log-likelihood of the Markov model is obtained numerically, while those of the semi-Markov model are estimated via Monte Carlo (Appendix A.6).
Figures 3(b) and 3(b) display the pointwise estimates and 95% confidence bands for cumulative incidence of symptomatic infection and all-comer infection, including asymptomatic infection, obtained using the best fit model. The observed incidence of symptomatic infections, which is continuously observed, is fully captured by the confidence bands, an indication that the semi-Markov model is a good fit. We observe a large increase in incidence of infections on day 7 (Figure 3(b)) — these are participants who were infected early, so are pre-/asymptomatic, and who have not yet been identified as infected. Thus, the model smoothly estimates a sharp increase in infections shortly after randomziation and avoids the artifactual bump at day 7 that is an artifact of weekly PCR assessment. The discrepancy between the observed and estimated incidence is greatest in the first week, when the majority of infections occurred — 64.5% (95% CI: 57.2%, 72.0%) of the total incidence in the placebo arm and 63.3% (95% CI: 55.6%, 71.3%) in the mAb arm. Furthermore the model has higher cumulative incidence at day 28, which reflects infections of short duration that are missed by intermittent sampling.
@@
Finally, we find that REGEN-COV reduces the mean duration of detectable viral shedding from 13.0 days (95% CI: 11.5 days, 14.6 days) to 6.2 days (95% CI: 5.0 days, 7.8 days). The estimated durations of PCR positivity are shorter than the estimates reported in O’Brien and others (2021) since we allow for seropositive participants to contribute information about the duration of viral shedding, even if they never develop symptoms or test positive by PCR. One obvious consequence of mAb participants shorter duration of PCR positivity is that the probability of detecting infection by PCR alone, i.e., without appealing to symptoms or serology and assuming that PCR assessments were conducted exactly weekly, is lower on the mAb arm; 0.69 (95% CI: 0.58, 0.78) versus 0.90 (95% CI: 0.85, 0.94). Thus, data assimilation is critical to capturing the full extent of infection since mAb reduces the sensitivity of symptoms and PCR for identifying cases.
@@
Figure 3: Cumulative incidence of disease (3(b)) and infection (3(b)) by arm, estimates of key functionals of interest, and confidence intervals based on 1,000 bootstrap samples. Solid lines correspond to the pointwise estimates, the shaded areas are $80\%$ and $95\%$ pointwise confidence band, and the dotted lines are the observed incidence. In Panel 3(c), the dots are the maximum likelihood estimates and the bars are $95\%$ confidence intervals.
@@ 5 Discussion
Our computational framework is closest, in spirit, to the Bayesian data augmentation MCMC algorithm of Barone and Tancredi (2022), who also employ a Markov surrogate to propose semi-Markov sample paths conditionally on the data. While data augmentation MCMC may be tractable for fitting simple models to small datasets, it is difficult to see this as a general solution in more complicated modeling scenarios when it becomes increasingly expensive to sample paths at each MCMC iteration. Furthermore, convergence and mixing deteriorate as the data become increasingly sparse relative to dynamics of the process (Papaspiliopoulos and others, 2007). These problems are compounded when poor convergence and mixing necessitate longer MCMC runs to obtain an adequate posterior sample. Nevertheless, we caution that our MCEM algorithm is not a panacea for non-identifiability despite its being agnostic about the structure of the model and accommodating a variety of measurement processes. General conditions under which MLEs for intermittently observed Markov processes may fail to exist, or are practically not identifiable, are characterized in Bladt and Sørensen (2005). The intuition from these results, such as needing sufficient temporal granularity in the data to ensure practical identifiability, applies equally to semi-Markov models. Scenarios that are computationally challenging for maximum likelihood estimation often prove difficult for Bayesian computation in the absence of well-calibrated priors. In our experience, our algorithm tends to fail “loudly” in settings where a responsible data analyst would reasonably expect that a model is over-parameterized given the extent of the data. We view obvious computational failure as a strength of the algorithm since it indicates that the model is not supported by the data.
@@ 6 Software
The methods presented in this manuscript are implemented in the MultistateModels.jl Julia package, available at https://github.com/fintzij/MultistateModels.jl. An ‘R‘ package making the Julia software callable from ‘R‘ is available at https://github.com/ammateja/multistatemodels together with vignettes demonstrating how to simulate from and fit various multistate models. A repository with archived versions of these packages and all simulation code that can be used to reproduce the results in this manuscript is available at https://github.com/ammateja/MultistateModelsPaper.
@@ 7 Supplementary Material
This article reflects the views of the author and should not be construed to represent FDA’s views or policies.
@@ Acknowledgments
The authors would like to thank to Michael P. Fay, Jing Wang, Arthur Winer, and Jennifer Gillman for their invaluable help that greatly improved this manuscript. This work utilized the computational resources of the NIH HPC Biowulf cluster (https://hpc.nih.gov). This project has been funded in whole or in part with federal funds from the National Cancer Institute, National Institutes of Health, under Contract No. 75N91019D00024. The content of this publication does not necessarily reflect the views or policies of the Department of Health and Human Services, nor does mention of trade names, commercial products, or organizations imply endorsement by the U.S. Government. Conflict of Interest: None declared.
@@ A.2 MCEM Q-function and importance weights
$Q\left(\boldsymbol{\theta},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\int l(\mathbf{Z},\mathbf{Y}\mid\boldsymbol{\theta})f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}\mathbf{Z}$
$=\int\left(\sum_{i=1}^{P}l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})\right)f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}\mathbf{Z}$
$=\sum_{i=1}^{P}\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}\mathbf{Z}$
$=\sum_{i=1}^{P}\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}Z_{i},$ (S1)
where $f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\widehat{\boldsymbol{\theta}}^{(t)}\right)=\prod_{i=1}^{P}f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ is the density of $\mathbf{Z}$ conditional on the data, and $f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ the density of $Z_{i}$ conditional on the data.
We approximate each integral, $\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)})\mathrm{d}Z_{i}$, in (A.2) using importance sampling. Let $h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)$ be another density with parameters $\boldsymbol{\theta}^{\prime}$ such that $f_{i}^{\prime}$ is absolutely continuous with respect to $h_{i}^{\prime}$. In our case, $f_{i}^{\prime}$ is the density of the semi-Markov process and $h_{i}^{\prime}$ is the density of the Markov surrogate, both conditional on the data. Rewriting the integral as
$\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}Z_{i}$ $=\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})\frac{f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)}h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)\mathrm{d}Z_{i},$
@@
$\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})\frac{f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)}h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)\mathrm{d}Z_{i}\approx\frac{1}{M_{i}}\sum_{m=1}^{M_{i}}\nu_{i,m}^{\prime(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\boldsymbol{\theta}\right).$ (S2)
@@
$\nu_{i,m}^{\prime(t-1)}$ $=\frac{f_{i}^{\prime}\left(Z_{i}^{(m)}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}^{\prime}\left(Z_{i}^{(m)}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)},$
$=\frac{f_{i}\left(Z_{i}^{(m)}\mid\widehat{\boldsymbol{\theta}}^{(t-1)}\right)p_{i}\left(\mathbf{Y}_{i}\mid Z_{i}^{(m)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)/g_{i}\left(\mathbf{Y}_{i}\mid\widehat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\boldsymbol{\theta}^{\prime}\right)q_{i}\left(\mathbf{Y}_{i}\mid Z_{i}^{(m)},\boldsymbol{\theta}^{\prime}\right)/r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)},$
$=\frac{f_{i}\left(Z_{i}^{(m)}\mid\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathbb{I}\left\{Z_{i}^{(m)}\text{ concords with }\mathbf{Y}_{i}\right\}/g_{i}\left(\mathbf{Y}_{i}\mid\widehat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\boldsymbol{\theta}^{\prime}\right)\mathbb{I}\left\{Z_{i}^{(m)}\text{ concords with }\mathbf{Y}_{i}\right\}/r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)},$
$\propto\frac{f_{i}\left(Z_{i}^{(m)}\mid\widehat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\boldsymbol{\theta}^{\prime}\right)}=\nu_{i,m}^{(t-1)}.$ (S3)
@@
$l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}Z_{i}\approx\frac{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)}{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}},$ (S4)
@@
$\widetilde{Q}\left(\boldsymbol{\theta},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\sum_{i=1}^{P}\frac{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)}{\sum_{j=1}^{M_{i}}\nu_{i,j}^{(t-1)}}.$

### v3

@@ 4.2 Model for clinical and immunological progression
We fit $6$ models, summarized in Tables S14 and S13, that were distinguished by the parameterization of their baseline transition intensities, i.e., $\eta_{k\ell}(\tau_{n},\boldsymbol{\gamma}_{k\ell})$ in equation (2). The most flexible model used piecewise linear baseline intensities with 3 interior knots at 3.5, 10.5, and 17.5 days for the transitions from naïve to PCR+ and from PCR+ to symptomatic, and 1 interior knot at 1 week for each of the PCR+ to PCR- transitions. A Markov model with time-homogeneous exponential intensities was the most rigid fully-parametric model. All intensities were adjusted for mAb assignment via a proportional intensity parameterization.
@@ 4.2.1 Seroconversion and asymptomatic infection
Let $C_{A}$, $C_{S}$ and $C=C_{A}+C_{S}$ denote the number of asymptomatic, symptomatic, and total seroconversions observed in an arm. Also let $\hat{I}_{A}$, $\hat{I}_{S}$, and $\hat{I}$ be the analogous number of infections estimated by the model. The estimated probabilities of seroconversion are $\hat{V}_{A}=C_{A}/\hat{I}_{A}$, $\hat{V}_{S}=C_{S}/\hat{I}_{S}$, and $\hat{V}=C/\hat{I}$. Note that we use the estimated probability of asyptomatic infection since otherwise $\hat{I}_{S}+\hat{I}_{A}$ may not equal $\hat{I}$, which would not be coherent under the model. Our use of the bootstrap is motivated by the observation that model-based asymptotic confidence intervals will understate the uncertainty in seroconversion following asymptomatic infection by virtue of conditioning on the number of asymptomatic seroconversion events.
@@ 4.3 Results
Table S14 compares the fitted models in terms of AIC. To showcase the improvement of the semi-Markov models over the Markov model, the difference in log-likelihood with the Markov model is displayed. The semi-Markov models offer a gain in log-likelihood between 78.7 and 101 units and the trade-off at the cost of additional model parameters is worthwhile, as reflected in the smaller AIC for more flexible models. The preferred model by AIC was parameterized with degree 1 B-splines with 2 interior knots for the $1\rightarrow 2$ and $2\rightarrow 4$ transitions, and 1 interior knot for the $2\rightarrow 3$ and $4\rightarrow 5$ transitions. The Markov model exponential baseline intensities had, by far, the worst AIC, while the model with Weibull intensities was worst among the semi-Markov models.
Figures 3(b) and 3(b) display the pointwise estimates and 95% confidence bands for cumulative incidence of symptomatic infection and all-comer infection, including asymptomatic infection, obtained using the best fit model. The observed incidence of symptomatic infections, which is continuously observed, is fully captured by the confidence bands, an indication that the semi-Markov model is a good fit. We observe a large increase in incidence of infections on day 7 (Figure 3(b)) — these are participants who were infected early, so are pre-/asymptomatic, and who have not yet been identified as infected. Thus, the model smoothly estimates a sharp increase in infections shortly after randomziation and avoids the artifactual bump at day 7 that is an artifact of weekly PCR assessment. The discrepancy between the observed and estimated incidence is greatest in the first week, when the majority of infections occurred — 64.5% (95% CI: 57.2%, 72.0%) of the total incidence in the placebo arm and 63.3% (95% CI: 55.6%, 71.3%) in the mAb arm. Furthermore, the model has higher cumulative incidence at day 28, which reflects infections of short duration that are missed by intermittent sampling.
@@
Finally, we find that REGEN-COV reduces the mean duration of detectable viral shedding from 13.0 days (95% CI: 11.5 days, 14.6 days) to 6.2 days (95% CI: 5.0 days, 7.8 days). The estimated durations of PCR positivity are shorter than the estimates reported in O’Brien and others (2021) since we allow for seropositive participants to contribute information about the duration of viral shedding, even if they never develop symptoms or test positive by PCR. One obvious consequence of mAb participants having a shorter duration of PCR positivity is that the probability of detecting infection by PCR alone, i.e., without appealing to symptoms or serology and assuming that PCR assessments were conducted exactly weekly, is lower on the mAb arm; 0.69 (95% CI: 0.58, 0.78) versus 0.90 (95% CI: 0.85, 0.94). Thus, data assimilation is critical to capturing the full extent of infection since mAb reduces the sensitivity of symptoms and PCR for identifying cases.
@@
Figure 3: Cumulative incidence of disease (3(b)) and infection (3(b)) by arm, estimates of key functionals of interest, and confidence intervals based on 1,000 bootstrap samples. Solid lines correspond to the pointwise estimates, the shaded areas are $80\%$ and $95\%$ pointwise confidence band, and the dotted lines are the observed incidence. In Panel 3(c), the dots are the maximum likelihood estimates and the bars are $95\%$ confidence intervals. Additional detailed results are reported in Table S15.
@@ 5 Discussion
Our computational framework is closest, in spirit, to the Bayesian data augmentation MCMC algorithm of Barone and Tancredi (2022), who also employ a Markov surrogate to propose semi-Markov sample paths conditionally on the data. While data augmentation MCMC may be tractable for fitting simple models to small datasets, it is difficult to see this as a general solution in more complicated modeling scenarios when it becomes increasingly expensive to sample paths at each MCMC iteration. Furthermore, convergence and mixing deteriorate as the data become increasingly sparse relative to dynamics of the process (Papaspiliopoulos and others, 2007). These problems are compounded when poor convergence and mixing necessitate longer MCMC runs to obtain an adequate posterior sample.
We caution that our MCEM algorithm is not a panacea for non-identifiability despite its being agnostic about the structure of the model and accommodating a variety of measurement processes. General conditions under which MLEs for intermittently observed Markov processes may fail to exist, or are practically not identifiable, are characterized in Bladt and Sørensen (2005). Their conclusions, such as needing sufficient temporal granularity in the data to ensure practical identifiability, apply equally to semi-Markov models. Scenarios that are computationally challenging for maximum likelihood estimation often prove difficult for Bayesian computation in the absence of well-calibrated priors. In our experience, our algorithm tends to fail “loudly” in settings where a responsible data analyst would reasonably expect that a model is over-parameterized given the extent of the data. We view obvious computational failure as a strength of the algorithm since it indicates that the model is not supported by the data.
@@ 6 Software
The methods presented in this manuscript are implemented in the MultistateModels.jl Julia package, available at https://github.com/fintzij/MultistateModels.jl. An ‘R‘ package making the Julia software callable from ‘R‘ is available at https://github.com/ammateja/multistatemodels together with vignettes demonstrating how to simulate from and fit various multistate models. Archived versions of these packages and code that can be used to reproduce this manuscript are available at https://github.com/ammateja/MultistateModelsPaper.
@@ 7 Supplementary Material
The content of this publication does not necessarily reflect the views or policies of the Department of Health and Human Services and the Food and Drug Administration, nor does mention of trade names, commercial products, or organizations imply endorsement by the U.S. Government.
@@ Acknowledgments
The authors would like to thank to Michael P. Fay, Jing Wang, Arthur Winer, and Jennifer Gillman for their invaluable help that greatly improved this manuscript. This work utilized the computational resources of the NIH HPC Biowulf cluster (https://hpc.nih.gov). This project has been funded in whole or in part with federal funds from the National Cancer Institute, National Institutes of Health, under Contract No. 75N91019D00024. Conflict of Interest: None declared.
@@ A.2 MCEM Q-function and importance weights
$Q\left(\boldsymbol{\theta},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\int l(\mathbf{Z},\mathbf{Y}\mid\boldsymbol{\theta})f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}\mathbf{Z}$
$=\int\left(\sum_{i=1}^{P}l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})\right)f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}\mathbf{Z}$
$=\sum_{i=1}^{P}\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}\mathbf{Z}$
$=\sum_{i=1}^{P}\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}Z_{i},$ (S1)
where $f^{\prime}\left(\mathbf{Z}\mid\mathbf{Y},\hat{\boldsymbol{\theta}}^{(t)}\right)=\prod_{i=1}^{P}f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ is the density of $\mathbf{Z}$ conditional on the data, and $f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ the density of $Z_{i}$ conditional on the data.
We approximate each integral, $\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)})\mathrm{d}Z_{i}$, in (A.2) using importance sampling. Let $h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)$ be another density with parameters $\boldsymbol{\theta}^{\prime}$ such that $f_{i}^{\prime}$ is absolutely continuous with respect to $h_{i}^{\prime}$. In our case, $f_{i}^{\prime}$ is the density of the semi-Markov process and $h_{i}^{\prime}$ is the density of the Markov surrogate, both conditional on the data. Rewriting the integral as
$\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}Z_{i}$ $=\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})\frac{f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)}h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)\mathrm{d}Z_{i},$
@@
$\int l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})\frac{f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)}h_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)\mathrm{d}Z_{i}\approx\frac{1}{M_{i}}\sum_{m=1}^{M_{i}}\nu_{i,m}^{\prime(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\boldsymbol{\theta}\right).$ (S2)
@@
$\nu_{i,m}^{\prime(t-1)}$ $=\frac{f_{i}^{\prime}\left(Z_{i}^{(m)}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}^{\prime}\left(Z_{i}^{(m)}\mid\mathbf{Y}_{i},\boldsymbol{\theta}^{\prime}\right)},$
$=\frac{f_{i}\left(Z_{i}^{(m)}\mid\hat{\boldsymbol{\theta}}^{(t-1)}\right)p_{i}\left(\mathbf{Y}_{i}\mid Z_{i}^{(m)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)/g_{i}\left(\mathbf{Y}_{i}\mid\hat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\boldsymbol{\theta}^{\prime}\right)q_{i}\left(\mathbf{Y}_{i}\mid Z_{i}^{(m)},\boldsymbol{\theta}^{\prime}\right)/r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)},$
$=\frac{f_{i}\left(Z_{i}^{(m)}\mid\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathbb{I}\left\{Z_{i}^{(m)}\text{ concords with }\mathbf{Y}_{i}\right\}/g_{i}\left(\mathbf{Y}_{i}\mid\hat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\boldsymbol{\theta}^{\prime}\right)\mathbb{I}\left\{Z_{i}^{(m)}\text{ concords with }\mathbf{Y}_{i}\right\}/r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)},$
$\propto\frac{f_{i}\left(Z_{i}^{(m)}\mid\hat{\boldsymbol{\theta}}^{(t-1)}\right)}{h_{i}\left(Z_{i}^{(m)}\mid\boldsymbol{\theta}^{\prime}\right)}=\nu_{i,m}^{(t-1)}.$ (S3)
@@
$l_{i}(Z_{i},\mathbf{Y}_{i}\mid\boldsymbol{\theta})f_{i}^{\prime}\left(Z_{i}\mid\mathbf{Y}_{i},\hat{\boldsymbol{\theta}}^{(t-1)}\right)\mathrm{d}Z_{i}\approx\frac{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)}{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}},$ (S4)
@@
$\tilde{Q}\left(\boldsymbol{\theta},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ $=\sum_{i=1}^{P}\frac{\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t-1)}l_{i}\left(Z_{i}^{(m)},\mathbf{Y}_{i}\mid\mathbf{X}_{i},\boldsymbol{\theta}\right)}{\sum_{j=1}^{M_{i}}\nu_{i,j}^{(t-1)}}.$